At any point of the path x=3cos⁡t,y=3sin⁡t,z=4t, what is the Normal vector


1
Expert's answer
2021-05-25T18:01:13-0400
"\vec r(t)=⟨3\cos t,3\sin t, 4t ⟩"

"\vec r\ '(t)=⟨-3\sin t,3\cos t, 4 ⟩"

"||\vec r\ '(t)||=\sqrt{(-3\sin t)^2+(-3\cos t)^2+(4)^2}=5"

"\vec T(t)=\dfrac{\vec r\ '(t)}{||\vec r\ '(t)||}=⟨-\dfrac{3}{5}\sin t,\dfrac{3}{5}\cos t, \dfrac{4}{5} ⟩"

"\vec T\ '(t)=⟨-\dfrac{3}{5}\cos t,-\dfrac{3}{5}\sin t, 0 ⟩"

"||\vec T\ '(t)||=\sqrt{(-\dfrac{3}{5}\cos t)^2+(-\dfrac{3}{5}\sin t)^2+(0)^2}=\dfrac{3}{5}"

"\vec N(t)=\dfrac{\vec T\ '(t)}{||\vec T\ '(t)||}=⟨-\cos t,-\sin t, 0 ⟩"